OpenStudy (anonymous):

Which of the following is the slope between the points (1, 5) and (5, 2)?

OpenStudy (anonymous):

Slope is -3/4

OpenStudy (anonymous):

m = (y2 -y1) / (x2 - x1) y2 = 2 , y1 = 5 x2 = 5 , x1 = 1 m = (2 - 5) / (5 - 1) m = -3 / 4 m is the slope

OpenStudy (anonymous):

@adjaharrison is right. I got confused with the formula sorry:)